OpenStudy (anonymous):

Consider the graph of the cosine function shown below. y= 4cos(2x) a. Find the period and amplitude of the cosine function. b. At what values of for do the maximum value(s), minimum values(s), and zeros occur?

OpenStudy (tkhunny):

What are the period and amplitude of \(y = \cos(x)\)

OpenStudy (anonymous):

the period of this equation is pie and the amplitude is 4

OpenStudy (anonymous):

i just dont know how to get the answers to part b

OpenStudy (scorcher219396):

The maximum and minimum of a normal cosine function are 1 (where the amplitude is 1), and since you have an amplitude of 4, you know that the y values for the maxes and mins will be 4. The period of a normal cosine function is 2pi, and you should know where those zeros occur (pi/2, 3pi/2). This function is half as long (period is only pi) so each of the zeros is half as far, meaning they're at pi/4 and 3pi/4. Same applies to find maxes and mins, the maxes of normal cosine are at 0 and 2pi, min at pi. These are all half as far, so the maxes are at 0 and pi, and the min is at pi/2. Your maxes are (0,4), (2pi, 4) and min at (pi/2, -4)

OpenStudy (tkhunny):

1) "pie" is for eating and "pi" is for mathematics an greek spellers. 2) Develop systematic thinking. \(y = \cos(x)\) Period: \(2\pi\) Amplitude: 1 Zeros: \(x = \pi/2+k\pi\) Min: \(x = \pi + 2k\pi\) Max: \(x = 0 + 2k\pi\) \(y = 4\cos(x)\) Period: \(2\pi\) Amplitude: 4 Zeros: \(x = \pi/2+k\pi\) Min: \(x = \pi + 2k\pi\) Max: \(x = 0 + 2k\pi\) \(y = 4\cos(2x)\) Period: \(\pi\) Amplitude: 4 Zeros: \(x = \pi/4+k\pi/2\) Min: \(x = \pi/2 + k\pi\) Max: \(x = 0 + k\pi\) Don't be afraid to write stuff out!